The age of Mary and Claire are in the ratio 5:7. Four years later their age will be in the ratio 3:4. Find their ages . please help me with this sum i am kind of getting confused and thanks!
Answer by stanbon(75887)

The age of Mary and Claire are in the ratio 5:7. Four years later their age will be in the ratio 3:4. Find their ages
-------------------
NOW DATA:
Mary is 5x yrs. old ; Claire is 7x yrs old.
---------------
4 years later DATA:
Mary is 5x+4 yrs old ; Claire is 7x+4 yrs old
----------------
EQUATION:
(5x+4)/(7x+4) = 3/4
Cross-multiply to get:
4(5x+4) = 3(7x+4)
20x+16 = 21x+12
x = 4
----------------
Ages NOW:
Mary = 5x = 20 yrs old
Claire = 7x = 28 yrs old
===========================
Cheers,
Stan H.